OpenStudy (anonymous):

A&X Electrical makes a lamp. The Research department found that the life of the lamp is normally distributed with a mean of 5000 hours and a standard deviation of 450 hours. 1. Determine the z-score for a lamp that lasts 4000 hours. z = 2.22 and the corresponding z score is .9865?

OpenStudy (anonymous):

those are my answers i just want to make sure that someone can confirm that they are correct

ganeshie8 (ganeshie8):

zscore has to be negative, since the observation is less than mean

OpenStudy (anonymous):

but the z is correct, right?

OpenStudy (anonymous):

5000-4000/450

OpenStudy (anonymous):

=2.22

jimthompson5910 (jim_thompson5910):

it's z = (x - mu)/sigma NOT z = (mu - x)/sigma

ganeshie8 (ganeshie8):

z = observation - mean / standard deviation

ganeshie8 (ganeshie8):

not the other way around

OpenStudy (anonymous):

x=5000, mu=4000

jimthompson5910 (jim_thompson5910):

the mean is 5000, so mu = 5000

OpenStudy (anonymous):

= -2.2

OpenStudy (anonymous):

z score = .0102?

jimthompson5910 (jim_thompson5910):

the z-score is -2.2 (roughly)

jimthompson5910 (jim_thompson5910):

I'm not sure what 0.0102 is referring to

OpenStudy (anonymous):

if you look at the table of z scores

jimthompson5910 (jim_thompson5910):

oh are you looking for the area under the curve to the left of z = -2.2 ?

OpenStudy (anonymous):

am not sure. all i have to do is answer the quesiton that i posted. so if im just supposed to find the corresponding z score and -2.2 is it then no

jimthompson5910 (jim_thompson5910):

I think they just want the z-score

jimthompson5910 (jim_thompson5910):

so z = -2.222... (the 2's after the decimal repeat) or just z = -2.22 or something like that should do

OpenStudy (anonymous):

ok thank you

jimthompson5910 (jim_thompson5910):

you're welcome

ganeshie8 (ganeshie8):

if u want the area under curve : http://www.wolframalpha.com/input/?i=zscore+-2.22

ganeshie8 (ganeshie8):

look at p-value

ganeshie8 (ganeshie8):

you should say : z = 2.22 and the corresponding area is .9865?

ganeshie8 (ganeshie8):

or : z = 2.22 and the corresponding probability is .9865?

jimthompson5910 (jim_thompson5910):

I think they just want the z-score ganeshie8

OpenStudy (anonymous):

okay thank you